OpenStudy (anonymous):

for a circle of radius 9 feet find the arc length subtended by a central angle of 52 degrees.

OpenStudy (ddcamp):

Finding the length of the arc?

OpenStudy (anonymous):

8.16

OpenStudy (anonymous):

right?

OpenStudy (ddcamp):

Hmm, I got 7.7. What steps did you take?

OpenStudy (anonymous):

im not sure

OpenStudy (ddcamp):

Steps I took: 1: Find the entire circumference C=pi*D=2*pi*r = 56.548 2: Find the fraction covered: There are 360° in a circle, and the arc substends 52°, which would be (52/360) times the entire circumference 3: Multiply 56.548*(52/360) = 8.168 You we're right, guess I typed it in wrong the first time. Sorry.

OpenStudy (anonymous):

yea thats what i did

OpenStudy (anonymous):

then what do i do?

OpenStudy (ddcamp):

8.168 is the answer for the length of that section. Is that what you're looking for?

OpenStudy (anonymous):

well my options are 13/5 pi, 11/5 pi, 13/11 pi or 5/11pi

OpenStudy (ddcamp):

Ah, so what we do is go back to our circimference and leave it in terms of pi: C=18(pi) Now, multiply that by 52/360 to get 13/5 pi

OpenStudy (anonymous):

oooh ok thank you !